CSePub Home Book Store Instructors Independent Authors About ANmrepots net icorneof$477$1500for 2013 TheMmhas a dividend papout ratio or2200% Themnanenty has $96382500 in debt and $1.865,400 00 Te trm pays 600% ania merestonther otstann att shareholder equity Answer Format Cumency Round to: 0 decimal places Entet Aswer Hore
Answer #1

Present debt equity ratio = debt/equity

=963825/1865400

=51.6685%

Next year equity = previous equity+ retained earnings

= 1865400+477875*(1-0.23)

=2233364

Maximum debt as per debt equity ratio = 2233364*51.6685% = 1153946

Debt issued = 1153946- 963825= 190121